OpenStudy (anonymous):

Is there a method to use besides trig sub for this problem? If not, How can I solve it with T.S.?

OpenStudy (anonymous):

\[\int\limits_{}^{}1/(3x ^{2}-81)\]

OpenStudy (anonymous):

it doesn't factor

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

partial fraction won't work

OpenStudy (anonymous):

3(x^2-27)

OpenStudy (anonymous):

but that does not help

OpenStudy (anonymous):

Yea, I did that to get \[1/3\int\limits_{}^{}1/(x ^{2}-27)\]

OpenStudy (anonymous):

The way I do trig sub is to make a triangle

OpenStudy (anonymous):

Some trig sub problem can be done easily instead with polar coordinate but not this one

OpenStudy (anonymous):

And in this scenario the hyp is x, adjacent is \[\sqrt(27)\] and the opposite is \[\sqrt(x^2-27)\] ?

OpenStudy (anonymous):

Not 100% positive about those sides

OpenStudy (anonymous):

why not partial fractions? \[\frac{1}{3}\int \frac{dx}{(x+3\sqrt{3})(x-3\sqrt{3})}\]

OpenStudy (anonymous):

Yea that would work but I wouldn't have seen it on a test lol

OpenStudy (anonymous):

can't imagine another way though.

OpenStudy (anonymous):

Thanks, how about trig sub approach?

OpenStudy (anonymous):

maybe hyperbolic? look in a book for that

OpenStudy (anonymous):

I think we might sub Sec(t) for x

OpenStudy (anonymous):

Yea I think thats what wolfram did there but I haven't even learned hyperbolic yet

OpenStudy (anonymous):

yeah fine but look at the answer. you get log something plus log something so i am thinking partial fractions is the way to go

OpenStudy (anonymous):

yay, partial

OpenStudy (anonymous):

Okie dokes

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

not that big of a deal i think. write \[(x+3\sqrt{3})A+(x-3\sqrt{3})B=1\]

OpenStudy (anonymous):

what about the 1/3?

OpenStudy (anonymous):

Does it just hang out till we get the answer?

OpenStudy (anonymous):

the integration*

OpenStudy (anonymous):

put \[x=3\sqrt{3}\] get \[ 6\sqrt{3}A=1\] \[A=\frac{1}{6\sqrt{3}}\] etc.

OpenStudy (anonymous):

yeah just leave the 1/3 outside

OpenStudy (anonymous):

and don't get married to wolfram's answer since there are many ways to write a log

OpenStudy (anonymous):

ok lemme take a swing at this with partial, thanks

OpenStudy (anonymous):

i will be without looking that if \[A=\frac{1}{6\sqrt{3}}\] then \[B=-\frac{1}{6\sqrt{3}}\] t

OpenStudy (anonymous):

enjoy